# Fleming-Viot helps speed up variational quantum algorithms in the presence of barren plateaus

### v2

@@ Abstract
Inspired by the Fleming-Viot stochastic process, we propose a parallel implementation of variational quantum algorithms with the aim of helping the algorithm get out of barren plateaus, where optimization direction is unclear. In the Fleming-Viot tradition, parallel searches are called particles. In our proposed approach, the search by a Fleming-Viot particle is stopped when it encounters a region where the gradient is too small or noisy, suggesting a barren plateau area. The stopped particle continues the search after being regenerated at another location of the parameter space, potentially taking the exploration away from barren plateaus. We first analyze the behavior of the Fleming-Viot particles from a theoretical standpoint. We show that, when simulated annealing optimizers are used as particles, the Fleming-Viot system is expected to find the global optimum faster than a single simulated annealing optimizer, with a relative efficiency that increases proportionally to the percentage of barren plateaus in the domain. This result is backed up by numerical experiments carried out on synthetic problems as well as on instances of the Max-Cut problem, which show that our method performs better than plain simulated annealing when large barren plateaus are present in the domain.
1 Introduction
@@
In both academia and industry, there has been considerable interest in solving optimization problems [1] using near-term quantum computers [2], albeit often in the form of quantum annealers [3, 4, 5, 6, 7, 8, 9, e.g.], or in the form of variational quantum algorithms (VQAs) on gate-based quantum computers [10, 11, 12, 13, 14, e.g.] or some combination of the two [15, e.g.]. VQAs [11, 13] are sometimes suggested as potential candidates to demonstrate an early quantum advantage in the near-term noisy quantum computers due to their simplicity, robustness to noise, and because the resources required are shared between a quantum computer and a classical optimizer. So far, VQAs have shown some potential in applications in a number of domains, including chemistry [16, 17, 18, 19, 20, 21], strongly correlated systems in condensed matter physics [22, 23, 24, 25, 26], combinatorial optimization [10, 27, 28, 29, 30, 31, 32, 33, 34], solving linear and nonlinear equations [35, 36, 37], supervised quantum machine learning [38, 39], generative models [40, 41, 42], and quantum neural networks [43, 44, 45, 46], to name a few.
Figure 1: The VQA schema. In dotted square the classical-quantum feedback loop. The circuit is initialized with parameters $\vartheta$ with the purpose of obtaining the empirical estimate of the expectation value of a set of observables $\{O_{k}\}_{k\in\mathbb{N}}$. In the classical part, a classical optimizer minimizes the cost function in order to obtain a new set of parameters, $\vartheta^{*}$ to be fed back to the quantum circuit.
@@
Variational quantum algorithms work by first running a parameterized quantum circuit that encodes the problem of interest, measuring the expected value of some observable relevant to this problem, classically learning the parameters that minimize a cost function related to the expected value of the observable, and repeating until the process converges to some satisfactory value. See Fig. 1.
1.1 Challenges encountered in VQAs
@@
Despite the potential of variational quantum algorithms, there exist several challenges in their deployment to practical problems. Concretely, the optimization problems involved are usually highly non-convex, and in general it is expected that they require super-exponential time (in the number of variables) to find the global optimum. Furthermore, in their practical implementation, the classical optimization subroutine is done with gradient-based algorithms for which in general there is no guarantee that such algorithms will converge to a global minimum in a reasonable amount of time [47, 48, 49, 50, 51]. The algorithm may get trapped in (suboptimal) local minima or saddle points. The iteration complexity as well as the impact of the biased noise present in NISQ [^1] Noise Intermediate-Scale Quantum devices are studied in [52]. Further convergence results are presented in [53].
@@
One major drawback of variational quantum algorithms is that they suffer from the so-called barren plateau (see Def. 1) problem [54, 55, 56, 57, 58, 59] especially when the problem is approached with hardware efficient ansätze. Barren plateaus are large regions in the parameter space where the gradient of the cost function vanishes or is noisy around zero, which, in turn, makes efficient optimization very difficult. Concretely, most gradient descent algorithms work as
@@
1.2 Our contribution
The usage of multiple parallel random walks to expedite coverage [78] can inspire similar potentially heuristic strategies in this context and to some extent our proposal follows along this line. As we will explain in detail in Sec. 2, by employing multiple parallel gradient descent initializations with a specific killing and regeneration rule, the so-called Fleming-Viot inspired algorithm, described in Alg. 1 of App. A, seems to better navigate the parameter landscape, efficiently circumventing barren plateaus by taking the paths towards other locations that potentially avoid them.
The rest of the paper is organized as follows: in Sec. 2 we introduce the Fleming-Viot process and describe how we can apply its principles to the classical optimization step of VQAs, which gives rise to Alg. 1. In Sec. 3 we provide and prove a theorem on the reduced expected hitting time of our method against vanilla simulated annealing. In Sec. 4 we showcase an application of our Fleming-Viot inspired algorithm for a synthetic problem with varying percentages of barren plateaus in the parameter landscape, as well as an application to a Quantum Approximation Optimization Algorithm (QAOA, a special case of VQA) for an instance of the Max-Cut problem on a graph with 8 nodes. We summarize and conclude in Sec. 5.
2 Fleming-Viot Process within VQAs
@@
3 Theoretical Analysis
Our main result is a hitting-time analysis for Alg. 1. In the next section (Sec. 4), the analysis is corroborated by numerical results. To present the analysis, let us define several quantities we will refer to:
@@ 3 Theoretical Analysis
Definition 1 (Barrren plateau of a function $f$)
@@
Assumption 1 (Stationary points and regions)
@@
Assumption 2 (Behaviour in an $\alpha$-barren plateau)
@@
Theorem 1 (Hitting-time analysis)
@@
Proof. Consider a process $X_{k}$ driven by SA with FV and $\hat{X}_{k}$ driven by SA without it. Consider the beginning of an event as the initialization of SA, at $k=0$, at some random non-barren point.
With $\alpha$ the gradient threshold and a decreasing learning rate $\eta_{k}\sim k^{-1/2}$, the hitting time to a stationary point is $C_{s}D\alpha^{-4}$ [83, Theorem 3.3]. We consider that generally there is an equal probability of the iterate hitting any one of the $s$ stationary points. Let us consider each type of point and its consequences on the two algorithms’ iterations.
@@
Otherwise, with probability $a/s$, both algorithms take $C_{a}\alpha^{-1}$ [83, Lemma 4.2] to escape a saddle point. Subsequently the next iterated event begins, i.e., we have another SA path from an agnostic starting point proceeding until it hits a stationary point.
Alternatively, with probability $(l-1)/s$, both algorithms take $C_{l}\alpha^{-1}$ [84, Corollary 6] to escape a local minimum. Subsequently, the next iterated event begins.
@@
The expected number of stationary points traversed before a global minimizer is reached is a geometric random variable with parameter $1/s$. Thus, for an expected $s-1$ time steps, we hit the other stationary points, and this enables us to compute the expected hitting time as the $s-1$ sum of the probabilities of hitting the different types of stationary points and their expected contribution to the total iteration time.
@@
Figure 2: Illustration of the FV optimization algorithm using $N=3$ particles, showing the two options described in Sec. 2 of particle regeneration after it is considered absorbed or killed: (i) particle reactivation (a.k.a. exploitation strategy), where an absorbed particle is regenerated to the position of one of the other particles selected uniformly at random, and (ii) particle reinitialization (a.k.a. exploration strategy), where an absorbed particle is regenerated to a randomly selected point in the parameter space, according to some probability distribution (e.g. uniform distribution). In the picture, particle $p_{1}$ is never absorbed, while particles $p_{2}$ and $p_{3}$ are absorbed at two different barren plateaus, for which both a reactivation and reinitialization paths are shown.
4 Numerical Experiments
@@
Two exploration rates are considered for the FV optimizer in order to analyze the effect of the exploitation vs. exploration strategies, which are controlled precisely by the exploration rate parameter. Note that, in the $100\%$ exploration rate, all killed particles are regenerated under the exploration strategy (described in Sec. 2).
@@
The FV parameters that are common to both problems are listed in Tab. 1.
@@
Table 1: FV parameters used in both the synthetic and Max-Cut experiments, as defined in Alg. 1 of App. A.
@@
Each set of experiments is defined by a parameterization specific to the optimization problem under consideration, described in the following two sections. Each experiment was run for $T=50$ optimization steps per particle, and the results are jointly presented in Sec. 4.3.
4.1 Synthetic Experiments
@@
With the goal of understanding the behaviour of the optimization methods on different barren levels, synthetic functions are generated using three different desired nominal barren percentages: 25%, 50%, 80%, aiming at generating function instances with low, medium and high barren levels. The precise definition of these three barren levels is given in Section 4.3. Each accepted [^2] As described in App. B. synthetic function is classified into one of these barren levels based on the actual percentage of barren plateaus present in the function, estimated as described in App. C. Finally, in order to obtain a balanced distribution of functions assigned to all three barren levels, synthetic functions are generated until 9 functions are accepted for analysis in each level.
@@
Tab. 2 lists the parameters defining the synthetic optimization problems.
@@
Table 2: Parameters defining the synthetic problems.
4.2 QAOA Experiments on the Max-Cut problem
@@
The experiments were carried out on the QAOA approximation of the Max-Cut problem for four different connectedness levels: $25\%,50\%,75\%,100\%$ on an 8-node graph with edge weights chosen uniformly at random between 0 and 1. The number of shots for the evaluation of the quantum circuit is 512. Tab. 3 lists the parameters defining the Max-Cut optimization problems.
@@
Table 3: Parameters defining the Max-Cut problems.
4.3 Experimental results
@@
In order to make results comparable across the two problems and across the different experiments, we define a performance metric that is independent of the global minimum value attained by the function and of the function range. The performance metric is thus defined as the minimum value found by the analyzed method on the function after standardization to the interval $[0,1]$, as: $(f-f_{min})/(f_{max}-f_{min})$, where $f_{min}$ and $f_{max}$ are respectively the true global minimum and maximum of the realized function, computed on a regular grid of $100\times 100$ in the synthetic case, and of $50\times 50$ in the Max-Cut case [^3] Even though a natural performance metric would be the relative error between the minimum value found by the analyzed method and the true global minimum, it is a metric that is highly dependent on the global minimum value and the function range. For instance, a function with global minimum equal to 1.0 and global maximum equal to 1.5 will most likely allow the optimization methods reach a smaller relative error than the one obtained on a function with the same global minimum but a global maximum equal to 10. In fact, the maximum possible error for the latter function is 900%, whereas it is just 50% for the former function. Therefore, the relative error metric is not comparable across these two functions.. The best possible performance is thus 0 and the worst is 1444In certain cases, the best performance could be slightly negative, due to the fact that the true global minimum is computed on a regular grid, whereas the estimated global minimum is not restricted to the grid..
The performance metric is analyzed in terms of the amount of barren plateaus present in the optimization function, estimated using the procedure described in App. C as a percentage of the domain size. Note that, although we may think that in the synthetic case the measure of barrenness could be given by the nominal barren percentage specified in the synthesis procedure, this is not fully appropriate because the actual barren percentage (as visually observed in the realized function’s landscape) may be substantially different from the nominal one, due to the final smoothing step in the synthesis process  [^5] See the end of App. B for a restriction imposed on the relationship between the estimated barren percentage and the nominal one to accept the generated function as a valid synthetic function to be used in experiments.. For the final analysis of performance, the estimated barren percentages are used to group the generated functions into one of the following three barren levels: LOW [0%, 33%], MEDIUM (33%, 66%], and HIGH (66%, 100%), which are analyzed separately.
@@
Figure 5: Speed by estimated barren level with which each method reaches the estimated function minimum on the function instances standardized to the $[0,1]$ interval. (a) Synthetic case with 27 instances, (b) Max-Cut problem with 24 instances. Each panel shows the distribution of the median, computed as described in Section 4.3. The thick coloured curves represent the median of the distribution, and the bands span the distribution’s min-max range. Lower positions of the curves are better. Although 500 optimization steps are used in the “SA-1” method, only the first 50 are shown.
@@ Appendix A Fleming-Viot optimization algorithm
13:procedure OptimizeAndRecordGradient
14:     for $j=1$ to $N$ do
15:         Perform optimization step for particle $j$
16:         Record observed gradient for particle $j$
17:     end for
18:end procedure
19:procedure CheckAbsorptionAndRegenerateParticles
20:     for $j=1$ to $N$ do
21:         if avg gradient of particle $j$ over last $W$ steps < ref_gradient then
22:              $\text{rnd}\sim\mathcal{U}[0,1]$
23:              if $\text{rnd}<\epsilon$ then
24:                  ReinitializeParticle(j)
25:              else
26:                  ReactivateParticle(j)
27:              end if
28:         end if
29:     end for
30:end procedure
31:procedure ReinitializeParticle(j)
32:     Regenerate particle $j$ to a randomly chosen point in the parameter space
33:     Reset the history of recorded gradient values for particle $j$
34:end procedure
35:procedure ReactivateParticle(j)
36:     Regenerate particle $j$ to a randomly chosen particle ($r$) among non-killed particles
37:     Assign the last $W$ gradient values of particle $r$ as the last $W$ gradient values of particle $j$
38:     Reset the learning rate of particle $j$ to the initial learning rate of particle $r$ with the goal of increasing the divergence between the two particles, thus boosting exploration
39:end procedure

### v3

# Parallel variational quantum algorithms with gradient-informed restart to speed up optimisation in the presence of barren plateaus

@@ Abstract
Inspired by the Fleming-Viot stochastic process, we propose a parallel implementation of variational quantum algorithms with the aim of reducing the time spent by the algorithm in barren plateaus, where optimization direction is unclear. In the Fleming-Viot tradition, parallel searches are called particles. In the proposed approach, the search by a Fleming-Viot particle is stopped when it encounters a region where the gradient is too small or noisy, suggesting a barren plateau area. The stopped particle continues the search after being regenerated at another location of the parameter space, potentially taking the exploration away from barren plateaus. We first analyze the behavior of the Fleming-Viot particles from a theoretical standpoint. We show that, when simulated annealing optimizers are used as particles, the Fleming-Viot system is expected to find the global optimum faster than a single simulated annealing optimizer, with a relative efficiency that increases proportionally to the percentage of barren plateaus in the domain. This result is corroborated by numerical experiments carried out on synthetic problems as well as on instances of the Max-Cut problem, which show that our method performs better than plain simulated annealing when large barren plateaus are present in the domain.
I Introduction
@@
In both academia and industry, there has been considerable interest in solving optimization problems using near-term quantum computers [1, 2], often using variational quantum algorithms (VQAs) on gate-based quantum computers [3, 4, 5, 6, 7]. VQAs [4, 6] are sometimes suggested as potential candidates to demonstrate an early quantum advantage in the near-term noisy quantum (NISQ) computers due to their simplicity, robustness to noise, and because the resources required are shared between a quantum computer and a classical optimizer. So far, VQAs have shown some potential in applications in a number of domains, including chemistry [8, 9, 10, 11, 12, 13], strongly correlated systems in condensed matter physics [14, 15, 16, 17, 18], combinatorial optimization [3, 19, 20, 21, 22, 23, 24, 25, 26], solving linear and nonlinear equations [27, 28, 29], supervised quantum machine learning [30, 31], generative models [32, 33, 34], and quantum neural networks [35, 36, 37, 38], to name a few.
@@
VQAs are iterative algorithms, where each iteration runs a parameterized quantum circuit that encodes the problem of interest, measures the expected value of some observable relevant to this problem, and classically learns the parameters that minimize a cost function related to the expected value of the observable. These iterations are repeated until the process converges to some satisfactory value. See Fig. 1 overleaf for an illustratioon.
I-A Challenges encountered in VQAs
@@
Despite the potential of variational quantum algorithms, there exist several challenges in their deployment to practical problems. Concretely, the optimization problems involved in classically learning the parameters are non-convex, and it is expected that classical subroutines require super-exponential time in the number of variables to find the global optimum. In practical implementations, one often utilizes gradient-based classical optimization subroutines with no guarantee that such a subroutine will converge to a global minimum in a finite amount of time [39, 40, 41, 42, 43]. Specifically, the gradient-based algorithm may get trapped in (suboptimal) local minima or saddle points. The iteration complexity, as well as the impact of the biased noise present in NISQ devices, are studied in [44]. Further convergence results are presented in [45].
Figure 1: The VQA schema, as presented in [2]. In dotted square the classical-quantum feedback loop. The circuit is initialized with parameters $\vartheta$ with the purpose of obtaining the empirical estimate of the expectation value of a set of observables $\{O_{k}\}_{k\in\mathbb{N}}$. In the classical part, a classical optimizer minimizes the cost function in order to obtain a new set of parameters, $\vartheta^{*}$ to be fed back to the quantum circuit.
@@
One major drawback of variational quantum algorithms is that they suffer from the so-called barren plateau problem [46, 47, 48, 49, 50, 51], especially when the problem is approached with hardware efficient ansätze. Barren plateaus are large regions in the parameter space where the gradient of the cost function vanishes or is noisy around zero, which, in turn, makes efficient optimization very difficult. (See Def. 1 on page 1 for a formal definition.) Specifically, gradient descent iteration is:
@@
I-B Our contribution
The usage of multiple parallel random walks to expedite coverage [70] can inspire similar potentially heuristic strategies in this context and to some extent our proposal follows along this line. As we will explain in detail in Sec. II, by employing multiple parallel gradient descent initializations with a specific killing and regeneration rule, the so-called Fleming-Viot inspired algorithm, described in Alg. 1 of App. A, seems to better navigate the parameter landscape, efficiently circumventing barren plateaus by taking the paths towards other locations that potentially avoid them.
The rest of the paper is organized as follows: in Sec. II we introduce the Fleming-Viot process and describe how we can apply its principles to the classical optimization step of VQAs, which gives rise to Alg. 1. In Sec. III we provide and prove a theorem on the reduced expected hitting time of our method against vanilla simulated annealing. In Sec. IV we showcase an application of our Fleming-Viot inspired algorithm for a synthetic problem with varying percentages of barren plateaus in the parameter landscape, as well as an application to a Quantum Approximation Optimization Algorithm (QAOA, a special case of VQA) for an instance of the Max-Cut problem on a graph with 8 nodes. We summarize and conclude in Sec. V.
II Fleming-Viot Process within VQAs
@@
III Theoretical Analysis
Our main result is a hitting-time analysis for Alg. 1. In the next section (Sec. IV), the analysis is corroborated by numerical results. To present the analysis, let us define several quantities we will refer to:
@@ III Theoretical Analysis
Definition 1 (Barrren plateau of a function $f$).
@@
Assumption 1 (Stationary points and regions).
@@
Assumption 2 (Behaviour in an $\alpha$-barren plateau).
@@
Theorem 1 (Hitting-time analysis).
@@
Proof.
Consider a process $X_{k}$ driven by SA with FV and $\hat{X}_{k}$ driven by SA without it. Consider the beginning of an event as the initialization of SA, at $k=0$, at some random non-barren point.
With $\alpha$ the gradient threshold and a decreasing learning rate $\eta_{k}\sim k^{-1/2}$, the hitting time to a stationary point is $C_{s}D\alpha^{-4}$ [75, Theorem 3.3]. We consider that generally there is an equal probability of the iterate hitting any one of the $s$ stationary points. Let us consider each type of point and its consequences on the two algorithms’ iterations.
@@
Otherwise, with probability $a/s$, both algorithms take $C_{a}\alpha^{-1}$ [75, Lemma 4.2] to escape a saddle point. Subsequently the next iterated event begins, i.e., we have another SA path from an agnostic starting point proceeding until it hits a stationary point.
Alternatively, with probability $(l-1)/s$, both algorithms take $C_{l}\alpha^{-1}$ [76, Corollary 6] to escape a local minimum. Subsequently, the next iterated event begins.
@@
The expected number of stationary points traversed before a global minimizer is reached is a geometric random variable with parameter $1/s$. Thus, for an expected $s-1$ time steps, we hit the other stationary points, and this enables us to compute the expected hitting time as the $s-1$ sum of the probabilities of hitting the different types of stationary points and their expected contribution to the total iteration time. ∎
@@
Figure 2: Illustration of the FV optimization algorithm using $N=3$ particles, showing the two options described in Sec. II of particle regeneration after it is considered absorbed or killed: (i) particle reactivation (a.k.a. exploitation strategy), where an absorbed particle is regenerated to the position of one of the other particles selected uniformly at random, and (ii) particle reinitialization (a.k.a. exploration strategy), where an absorbed particle is regenerated to a randomly selected point in the parameter space, according to some probability distribution (e.g. uniform distribution). In the picture, particle $p_{1}$ is never absorbed, while particles $p_{2}$ and $p_{3}$ are absorbed at two different barren plateaus, for which both a reactivation and reinitialization paths are shown.
IV Numerical Experiments
@@
Two exploration rates are considered for the FV optimizer in order to analyze the effect of the exploitation vs. exploration strategies, which are controlled precisely by the exploration rate parameter. Note that, in the $100\%$ exploration rate, all killed particles are regenerated under the exploration strategy (described in Sec. II).
@@
The FV parameters that are common to both problems are listed in Tab. I.
@@
TABLE I: FV parameters used in both the synthetic and Max-Cut experiments, as defined in Alg. 1 of App. A.
@@
Each set of experiments is defined by a parameterization specific to the optimization problem under consideration, described in the following two sections. Each experiment was run for $T=50$ optimization steps per particle, and the results are jointly presented in Sec. IV-C.
IV-A Synthetic Experiments
@@
With the goal of understanding the behaviour of the optimization methods on different barren levels, synthetic functions are generated using three different desired nominal barren percentages: 25%, 50%, 80%, aiming at generating function instances with low, medium and high barren levels. The precise definition of these three barren levels is given in Section IV-C. Each accepted [^1] As described in App. B. synthetic function is classified into one of these barren levels based on the actual percentage of barren plateaus present in the function, estimated as described in App. C. Finally, in order to obtain a balanced distribution of functions assigned to all three barren levels, synthetic functions are generated until 9 functions are accepted for analysis in each level.
@@
Tab. II lists the parameters defining the synthetic optimization problems.
@@
TABLE II: Parameters defining the synthetic problems.
IV-B QAOA Experiments on the Max-Cut problem
@@
The experiments were carried out on the QAOA approximation of the Max-Cut problem for four different connectedness levels: $25\%,50\%,75\%,100\%$ on an 8-node graph with edge weights chosen uniformly at random between 0 and 1. The number of shots for the evaluation of the quantum circuit is 512. Tab. III lists the parameters defining the Max-Cut optimization problems.
@@
TABLE III: Parameters defining the Max-Cut problems.
IV-C Experimental results
@@
In order to make results comparable across the two problems and across the different experiments, we define a performance metric that is independent of the global minimum value attained by the function and of the function range. The performance metric is thus defined as the minimum value found by the analyzed method on the function after standardization to the interval $[0,1]$, as: $(f-f_{min})/(f_{max}-f_{min})$, where $f_{min}$ and $f_{max}$ are respectively the true global minimum and maximum of the realized function, computed on a regular grid of $100\times 100$ in the synthetic case, and of $50\times 50$ in the Max-Cut case  [^2] Even though a natural performance metric would be the relative error between the minimum value found by the analyzed method and the true global minimum, it is a metric that is highly dependent on the global minimum value and the function range. For instance, a function with global minimum equal to 1.0 and global maximum equal to 1.5 will most likely allow the optimization methods reach a smaller relative error than the one obtained on a function with the same global minimum but a global maximum equal to 10. In fact, the maximum possible error for the latter function is 900%, whereas it is just 50% for the former function. Therefore, the relative error metric is not comparable across these two functions.. The best possible performance is thus 0 and the worst is 1  [^3] In certain cases, the best performance could be slightly negative, due to the fact that the true global minimum is computed on a regular grid, whereas the estimated global minimum is not restricted to the grid..
The performance metric is analyzed in terms of the amount of barren plateaus present in the optimization function, estimated using the procedure described in App. C as a percentage of the domain size. Note that, despite setting the desired barren level in the synthetic function, the actual barren percentage of the realized function may be substantially different from the nominal value, due to the smoothing step of the function synthesis procedure  [^4] See the end of App. B for a restriction imposed on the relationship between the estimated barren percentage and the nominal one to accept the generated function as a valid synthetic function to be used in experiments.. For the final analysis of performance, the estimated barren percentages are used to group the generated functions into one of the following three barren levels: LOW [0%, 33%], MEDIUM (33%, 66%], and HIGH (66%, 100%), which are analyzed separately.
@@
Figure 5: Speed by estimated barren level with which each method reaches the estimated function minimum on the function instances standardized to the $[0,1]$ interval. (a) Synthetic case with 27 instances, (b) Max-Cut problem with 24 instances. Each panel shows the distribution of the median, computed as described in Section IV-C. The thick coloured curves represent the median of the distribution, and the bands span the distribution’s min-max range. Lower positions of the curves are better.
@@ Appendix A Fleming-Viot optimization algorithm
procedure OptimizeAndRecordGradient
     for $j=1$ to $N$ do
         Perform optimization step for particle $j$
         Record observed gradient for particle $j$
     end for
end procedure
procedure CheckAbsorptionAndRegenerateParticles
     for $j=1$ to $N$ do
         if avg gradient of particle $j$ over last $W$ steps ¡ ref_gradient then
              $\text{rnd}\sim\mathcal{U}[0,1]$
              if $\text{rnd}<\epsilon$ then
                  ReinitializeParticle(j)
              else
                  ReactivateParticle(j)
              end if
         end if
     end for
end procedure
procedure ReinitializeParticle(j)
     Regenerate particle $j$ to a randomly chosen point in the parameter space
     Reset the history of recorded gradient values for particle $j$
end procedure
procedure ReactivateParticle(j)
     Regenerate particle $j$ to a randomly chosen particle ($r$) among non-killed particles
     Assign the last $W$ gradient values of particle $r$ as the last $W$ gradient values of particle $j$
     Reset the learning rate of particle $j$ to the initial learning rate of particle $r$ with the goal of increasing the divergence between the two particles, thus boosting exploration
end procedure